OpenStudy (anonymous):

(-11/4+(-2/3) ^(-1)^2

zepdrix (zepdrix):

Mmmm is this kinda what the problem looks like?\[\Large \left[-\frac{11}{4}+\left(-\frac{2}{3}\right)^{-1}\right]^2\] Or is the -1 power being applied to both fractions?

OpenStudy (anonymous):